OpenStudy (anonymous):

What is the value of the expression 14 + 6 • 2 – 10 ÷ 5? A. 38 B. 6 C. 0 D. 24

OpenStudy (anonymous):

@inowalst

OpenStudy (anonymous):

@jeweaver6895

OpenStudy (anonymous):

i think b

OpenStudy (anonymous):

i am not stupid, i just want to be sure and see if u can do it

OpenStudy (anonymous):

i will give medal

OpenStudy (inowalst):

14 + 6 x 2-10 / 5 14+6=20 20x2=40 40-10=30 30/5=6

OpenStudy (inowalst):

You were right the first time.

OpenStudy (anonymous):

kk, thx

OpenStudy (inowalst):

You're welcome!